OpenStudy (anonymous):

wingspan owns a lawn mowing and landscaping business. The income from the business is given by the function f(x) = 2x + 54, where f(x) is the income in dollars and x is the area in square meters of lawn mowed. If wingspan has earned {204, 344, 450, 482} dollars in the last four months, what are the corresponding areas of lawn he mowed?

OpenStudy (anonymous):

someone please help me i have been trying this forever

OpenStudy (ikram002p):

its a leaner function you just have to find what is X in each cases case a : \[204=2x+54 \rightarrow x1= \frac{ 204-54 }{2 }\] case b :\[344=2x+54 \rightarrow x2= \frac{344-54 }{2 }\] case c :\[450=2x+54 \rightarrow x3= \frac{450-54 }{2 }\] case d :\[482=2x+54 \rightarrow x4= \frac{482-54 }{2 }\] I'll let you do the lest of work to find the corresponding areas of lawn he mowed which are represented by {x1,x2,x3,x4} good luck :)

OpenStudy (anonymous):

thanks

OpenStudy (ikram002p):

wlc :D

OpenStudy (anonymous):

i got a c thank u finally after 19 tries